OpenStudy (anonymous):

Six blocks of height 4h+4 each and 3 blocks of height 8-2h each are stacked on top of each other to form one big tower. Find an expression for the overall height of the tower

OpenStudy (adajiamcneal):

18h + 48

OpenStudy (triciaal):

total height = 6(4h + 4) + 3(8 - 2h) = 24(h + 1) + 6( 4 - h) (24 - 6) h + (24 + 24) 18 h + 48